# Deterministic and Statistical Analysis of the DoF of Continuous Linear Arrays in the Near Field

### v1

@@ V-C2 Full Visibility of the Rx Array
Figure 6: The DoF for $L_{T}=0.2m$, $L_{R}=2m$, $f=30GHz$, $y_{0}=0m$ and different normalized distance $x_{0}/L_{R}$.
@@
Figure 7: CCDF of the DoF versus $m_{th}$ for different values of the radius $R$ under partial and full visibility assumption and for $x_{0}$ being a random variable. Markers denote the analytical results.
Figure 8: Conditional CCDF of the DoF versus $m_{th}$ for different values of $L_{R}$ and $x_{0}$. Markers denote the analytical results.
@@
VI Numerical Results and Discussion
VI-A Deterministic Analysis
Because there are seven variables involved in the calculation of the number of DoF, namely $\theta_{T},\theta_{R},x_{0},y_{0},L_{T},L_{R},f$, the number of possible combinations is extremely large. Thus, in this subsection, representative cases are examined and numerical results are given. In all results presented, the frequency is set to $f=30\,GHz$, and the length of the transmitting surface is set to $L_{T}=0.2$ meters. In Fig. 5, the number of DoF for three different Rx locations is given assuming that a large receiving array of length $L_{R}=5$ meters is used. The number of DoF is given versus $\theta_{R}\in(-\pi/2,\pi/2]$, for various values of $\theta_{T}$. The line that connects the centers of the two arrays is the $x$-axis in the first plot, the $y$-axis in the second plot, and it forms an angle of $\pi/4$ with the $x$-axis in the third plot. The first observation is that a different maximum number of DoF is obtained for different values of $\theta_{T}$. In addition, different values are obtained for the maximum value and the minimum value of the DoF for different values of $\theta_{R}$. The second observation is the high number of modes available in the considered settings, for a small transmitting array and for a wide range of angles $\theta_{T}$ and $\theta_{R}$. An interesting conclusion based on the obtained results is that, if the location and the orientation of the surfaces are known to the large array, an efficient resource allocation strategy can be utilized, based on the number of DoF attainable at each small array. The impact of the distance between the two arrays on the number of DoF is depicted in Fig. 6, where various normalized to $L_{R}$ distances are examined. This result reveals the advantage provided by near-field communications. Depending on the distance between the two arrays and the rotation angle $\theta_{R}$, the number of DoF takes values in the range [0,40], as expected from (43). The zero value corresponds to the case where no visibility is achieved between the two arrays.
VI-B Statistical Analysis
In this subsection, the statistical behavior of the number of DoF is evaluated to gain system-level insights. The accuracy of the analytical results is verified against Monte-Carlo simulations. Unless stated otherwise, the following parameters are utilized: $f=30$ GHz, $L_{T}=0.2$ meters, $L_{R}=2$ meters, and $R=20$ meters. Fig. 7 shows the complementary cumulative distribution function (CCDF), $F^{c}_{m}(m_{th})$, of the number of DoF for different values of $R$ and for $x_{0}$ being a random variable. The CCDF $F^{c}_{m}(m_{th})$ is defined as $F^{c}_{m}(m_{th})=1-\mathbb{P}[m\leq m_{th}]$ and is obtained through numerical integration of Theorem 1 and Theorem 2 under the partial visibility assumption and Theorem 3 under the full visibility assumption. The first observation is that the probability of achieving a target value, e.g., $m_{th}=20$ modes, which corresponds to half the maximum number of DoF, $2L_{T}/\lambda$, dramatically increases with the decrease of $R$ in the full visibility scenario. This result i) highlights the need for establishing full visibility with the receiving array in mmWave networks, and ii) quantifies the available DoF as the links move from being in the near-field regime to being in the far-field regime. The Fraunhofer distance is $d_{FF}=2L_{R}^{2}/\lambda=800$ meters. Interestingly, the statistical behavior of the DoF is quite different between the partial and full visibility scenario for the smaller values of $R$ and the superiority in the number of DoF under full visibility is apparent. In fact, the probability that the number of DoF is greater than $m_{th}=20$ under the full visibility scenario for $R=5$ meters is slightly lower than 40%, while under the partial visibility scenario it is approximately 10%.
Fig. 8 shows the conditional $F^{c}_{m}(m_{th})$ versus $m_{th}$ for two values of the coordinate $x_{0}$ and for different values of $L_{R}$. The CCDF $F^{c}_{m}(m_{th})$ is obtained by numerical integration of Theorem 4. The first observation is that the maximum number of DoF is clearly location-dependent. Interestingly, a comb-like behavior is depicted for various values of $L_{R}$. Thus, the decrease in $F^{c}_{m}(m_{th})$ becomes steeper for larger receiving arrays and a higher number of DoF is more likely to be obtained. This is also a consequence of the higher probability that the receiving array is visible, which clearly depends on $x_{0}$ and $L_{R}$, as given by Definition 1.
@@ VII Conclusions
In this work, the number of DoF between a small linear array located in the near field of a large linear array is examined. The proposed deterministic framework is applicable to general network deployments beyond the widely studied paraxial setting. In order to derive the number of DoF, the framework is based on the calculation of mutual visibility conditions between the two arrays. To this end, apart from the length of the two arrays, four more parameters are considered, namely the Cartesian coordinates of the center of the receiving array and two angles that model the rotation of each array around their center. Subsequently, a stochastic geometry framework is proposed to capture the statistical behavior of the number of DoF. The analysis is performed for practical scenarios as a first step towards the investigation of the system-level performance in near-field communication networks. Among others, the results: i) quantified the advantage provided by near-field communications in mmWave networks, ii) highlighted the importance of the distance and relative orientation between the arrays, and iii) showed that, for the examined scenario, a preferred array association policy should be primarily based on the distance criterion and secondarily on the array length.

### v2

@@ VI-A Deterministic Analysis
(a) The DoF for $x_{0}=10m$, $y_{0}=0m$
(b) The DoF for $x_{0}=0m$, $y_{0}=10m$
(c) The DoF for $x_{0}=5m$, $y_{0}=5m$
Figure 8: Comparison of the number of DoF calculated using the proposed kernel-based and the SVD-based method, for $L_{T}=0.2m$, $L_{R}=5m$, $f=30\,$GHz.
The impact of the distance between the two arrays on the number of DoF is depicted in Fig. 9, where various normalized to $L_{R}$ distances are examined. This result reveals the advantage provided by near-field communications. Depending on the distance between the two arrays and the rotation angle $\theta_{R}$, the number of DoF takes values in the range [0,19], as expected from (47). The zero value corresponds to the case where no visibility is achieved between the two arrays.
Figure 9: The DoF for $L_{T}=0.2m$, $L_{R}=2m$, $f=30\,$GHz, $y_{0}=0m$, $\theta_{T}=0^{o}$, and different normalized distance $x_{0}/L_{R}$.
VI-B Statistical Analysis
In this subsection, the statistical behavior of the number of DoF is evaluated to gain system-level insights. The accuracy of the analytical results is verified against Monte-Carlo simulations. Unless stated otherwise, the following parameters are utilized: $f=30$ GHz, $L_{T}=0.2$ meters, $L_{R}=2$ meters, and $R=20$ meters. Fig. 10 shows the complementary cumulative distribution function (CCDF), $F^{c}_{m}(m_{th})$, of the number of DoF for different values of $R$ and for $x_{0}$ being a random variable. The CCDF $F^{c}_{m}(m_{th})$ is defined as $F^{c}_{m}(m_{th})=1-\mathbb{P}[m\leq m_{th}]$ and is obtained through numerical integration of Theorem 1 and Theorem 2 under the partial visibility assumption and Theorem 3 under the full visibility assumption. A key observation is that a decrease in the range $R$ of $x_{0}$ significantly increases the probability of having a specific number of DoF. Therefore, as the distance $x_{0}$ decreases, the relative orientation of the arrays dominates the resulting DoF. Furthermore, the probability of achieving a target value of, e.g., $m_{th}=20$ modes, which corresponds to approximately half of the maximum number of DoF, $2L_{T}/\lambda$, dramatically increases with the decrease of $R$ in the full visibility scenario. This result highlights i) the need for establishing full visibility with the receiving array, and ii) the fact that the available number of DoF decreases rapidly with distance $x_{0}$. The latter is observed in Fig. 10(b), where at a range of $R=200\,m$, i.e., well below the Fraunhofer distance $d_{FF}=968\,m$, the probability of achieving $m>1$ is very low. Interestingly, the statistical behavior of the DoF is quite different between the partial and full visibility scenario for the smaller values of $R$ and the superiority in the number of DoF in full visibility is apparent. In fact, the probability that the number of DoF is greater than $m_{th}=20$ in the full visibility scenario for $R=5$ meters is slightly lower than 40%, while in the partial visibility scenario it is approximately 10%.
@@ VI-B Statistical Analysis
Figure 10: CCDF of the DoF versus $m_{th}$ for different values of the radius $R$ under partial and full visibility assumption and for $x_{0}$ being a random variable. Markers denote analytical results and the solid lines Monte Carlo simulations.
Fig. 11 shows the conditional $F^{c}_{m}(m_{th})$ versus $m_{th}$ for two values of the coordinate $x_{0}$ and for different values of $L_{R}$. The CCDF $F^{c}_{m}(m_{th})$ is obtained by numerical integration of Theorem 4. The first observation is that the maximum number of DoF is clearly location-dependent. Interestingly, a comb-like behavior is depicted for various values of $L_{R}$. Thus, the decrease in $F^{c}_{m}(m_{th})$ becomes steeper for larger receiving arrays and a higher number of DoF is more likely to be obtained. This is also a consequence of the higher probability that the receiving array is visible, which clearly depends on $x_{0}$ and $L_{R}$, as given by Definition 1. Fig. 11 also shows that the increase in the number of DoF with the length of the Rx array $L_{R}$ is more profound for smaller values of $x_{0}$.
Figure 11: Conditional CCDF of the DoF versus $m_{th}$ for different values of $L_{R}$ and $x_{0}$. Markers denote analytical results and the lines Monte Carlo simulations.
@@ VII Conclusions
In this work, the number of DoF between a small and a large linear array is examined, with an emphasis on the near-field. The proposed deterministic framework is applicable to general network deployments beyond the widely studied paraxial setting. In order to derive the number of DoF, the framework is based on the calculation of mutual visibility conditions between the two arrays. To this end, apart from the length of the two arrays, four more parameters are considered, namely the Cartesian coordinates of the center of the receiving array and two angles that model the rotation of each array around their center. The results are shown to coincide with those obtained with the SVD-based method. Subsequently, a stochastic geometry framework is proposed to capture the statistical behavior of the number of DoF. The analysis is performed for practical scenarios as a first step towards the investigation of the system-level performance in near-field communication networks. Among others, the results: i) quantified the advantage provided by near-field communications in mmWave networks, ii) highlighted the importance of the distance and relative orientation between the arrays, and iii) showed that, for the examined scenario, a preferred array association policy should be primarily based on the distance criterion and secondarily on the array length. Two future research challenges have been identified: i) the investigation of the off-boresight near/far field boundaries for two large arrays in a wireless link, and ii) the extension of this work to planar arrays and an analysis in three dimensions.